OpenStudy (anonymous):

The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@mony01

OpenStudy (anonymous):

@shortycme

OpenStudy (anonymous):

hold on i have to figure out a way to explain how i get my answer so igreen doesn't get me suspended again

OpenStudy (mony01):

well the pythagorean theorem is a^2+b^2=c^2 right and if you do 6^2+8^2=10^2 you'll get 36+64=100 meaning whatever a^2+b^ is you'll get c^2. the same goes for the other problem

OpenStudy (cwrw238):

the numbers are the 3 sides of a right angles triangle

OpenStudy (anonymous):

Thanks bruh